How do you solve this quadratic equation? It doesn't appear in the standard quadratic form so I'm unsure how to solve the problems I have like this one problem. I do know how to solve the ones that look like f(x) = a x^2 + b x + c ; please help me with this:
x^3-2x^2+5x-6=0

Answer by mananth(16949)

use factor theorem.
f(x) = x^3+2x^2-5x-6
f(-1) = -1+2+5-6 = 0
So, x+1 is a factor

By synthetic division,
x^3+2x^2-5x-6 = (x+1)(x^2+x-6) = (x+1)(x-2)(x+3) = 0

x = -1, 2, -3
m.ananth@hotmail.ca